jason78687:

A cylinder with height 11 inches and radius 4 inches is filled with eggnog. It is dropped and begins to leak at a rate of 5 inches/sec how long will it take for height to reach 8 inches??

jason78687:

@jhonyy9

jhonyy9:

so given that the height is 11 inches and ask after how many time will take for height 8 inches 11-8 = 3 inches given that leak at a rate of 5 inches/sec using the rule of 3 simple we can write 5 inches in 1 sec 3 inches in x cross multiplie and calcule the x value x = 3×1/5 x = 3/5 sec x = 0,6 sec hope answered understandably easy

jason78687:

@jhonyy9 wrote:
so given that the height is 11 inches and ask after how many time will take for height 8 inches 11-8 = 3 inches given that leak at a rate of 5 inches/sec using the rule of 3 simple we can write 5 inches in 1 sec 3 inches in x cross multiplie and calcule the x value x = 3×1/5 x = 3/5 sec x = 0,6 sec hope answered understandably easy
Why not simply multiply 8 why measure from top?

jhonyy9:

@jason78687 wrote:
@jhonyy9 wrote:
so given that the height is 11 inches and ask after how many time will take for height 8 inches 11-8 = 3 inches given that leak at a rate of 5 inches/sec using the rule of 3 simple we can write 5 inches in 1 sec 3 inches in x cross multiplie and calcule the x value x = 3×1/5 x = 3/5 sec x = 0,6 sec hope answered understandably easy
Why not simply multiply 8 why measure from top?
please read the question of your posted problem

jhonyy9:

first the height is 11 inches and you need calcule when after how many sec will have the height 8 inches the difference is 3 inches

jason78687:

Well it says 5 inches fill in 1 sec i.e 1 inch in 1/5 sec So for 8 inches to be filled shouldnt it be 8/5 seconds

jhonyy9:

no i think you dont understand the question from the height of 11 inches ask when after how many sec will have height of 8 inches to get from 11 inches the 8 inches the difference is 3 inches so this 3 inches you need calcule in how many sec will fill ok ?

jason78687:

Oh okk got it.

jhonyy9: